OpenStudy (anonymous):

√(36x^4 ) simplify fullly

OpenStudy (anonymous):

hint.. \[\Huge \sqrt[a]{x^b}=x^{\frac ba} \]

Parth (parthkohli):

 => 36x * 36x^2 => 36x^3